# Infinite quantum signal processing for arbitrary Szegő functions

### v2

@@ 4.2. Riemann-Hilbert factorization: Proof of Theorem 5
where $(a_{-},b_{-})$ is the NLFT of a sequence supported on $(-\infty,-1]$ and $(a_{+},b_{+})$ is the NLFT of a sequence supported on $[0,\infty)$, and then it suffices to define
@@
$=\lim\limits_{\eta\to 0}\langle A,1\rangle=A(\infty)\,.$
@@ 4.3. Lipschitz estimate
$\left\|\begin{pmatrix}A\\ B\end{pmatrix}-\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\right\|=\sup\left|\langle\begin{pmatrix}C\\ D\end{pmatrix},\begin{pmatrix}A\\ B\end{pmatrix}-\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\rangle\right|\,,$
@@
$=\langle\begin{pmatrix}C\\ D\end{pmatrix},\left(\operatorname{Id}+M\right)^{-1}\left(M^{\prime}-M\right)\left(\operatorname{Id}+M^{\prime}\right)^{-1}\begin{pmatrix}1\\ 0\end{pmatrix}\rangle\,.$
@@
Because $\begin{pmatrix}C\\ D\end{pmatrix}\in\mathcal{F}$, then $(\operatorname{Id}-M)\begin{pmatrix}C\\ D\end{pmatrix}\in\mathcal{D}\cap\mathcal{D}^{\prime}$ and so by Eq. 60 we may write this last inner product as
@@
$=\langle\begin{pmatrix}0&(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})^{*}\\ -(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})&0\end{pmatrix}\left(\operatorname{Id}-M\right)^{-1}\begin{pmatrix}C\\ D\end{pmatrix},\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\rangle\,,$
@@
$=\langle\left(\operatorname{Id}-M\right)^{-1}\begin{pmatrix}C\\ D\end{pmatrix},\begin{pmatrix}0&-(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})^{*}\\ (\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})&0\end{pmatrix}\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\rangle\,,$
@@
uniformly in $n$. Without loss of generality, take $n$ to be 00. By [1, (6.13)] and then Eq. 80, we may write
@@
(87) $A(\infty)^{-1}\left\|B-B^{\prime}\right\|_{L^{2}(\mathbb{T})}+A(\infty)^{-1}A^{\prime}(\infty)^{-\frac{1}{2}}\left|\frac{B^{\prime}(0)}{(A^{\prime})(\infty)}\right|\left\|A-A^{\prime}\right\|_{L^{2}(\mathbb{T})}\,.$
@@ 4.4. Plancherel equality and maximal solution
We remark that the particular choice of $a(z)$ given by Theorem 4 actually corresponds to the maximal solution, which is a special class of solutions with symmetric phase-factor proposed in [31] and enjoys many desirable properties. To reveal the connection between the function $a(z)$ constructed by Theorem 4 and the maximal solution, we first review the construction of the $SU(2)$ matrix corresponding to the maximal solution. Without loss of generality, let $f(z)$ be a real even target polynomial of degree $2d$. Then we factorize
@@
is a polynomial and is in $H^{2}(\mathbb{D})$, and it satisfies $\left\lvert\widetilde{a}(z)\right\rvert^{2}=\mathfrak{F}(z)$. Following [31, Theorem 4], we construct
@@
which is the unitary matrix associated with the maximal solution. It is exactly the unitary matrix $U_{d}(x,\Psi)$ associated with $(a(z),b(z))$ by exploiting $a(z^{2})=z^{-4d}\widetilde{a}(z)$ and verifying Eq. 27. It also implies that the construction method we propose gives the maximal solution with theoretical guarantee, while many other algorithms don’t.
@@
If $(a,b)$ is the NLFT of a sequence $F\in\ell^{2}(\mathbb{Z})$, then by the inner-outer factorization theorem for $H^{\infty}(\mathbb{D})$ functions [12, Corollary 5.6, Chapter II], we may write $a^{*}=IO$ where $I$ and $O$ are inner and outer functions on $\mathbb{T}$, respectively. By Eq. 22 we may write
@@
where the last equality follows from Eq. 91. Thus as in [30, (3.1) and the Remark on p.16], we have the nonlinear Plancherel identity
@@ 4.5. Proof of Theorem 1
A careful readthrough of this section up till now, along with the existence and uniqueness proofs above yields the following, which is a more comprehensive version of Lemma 6.
@@
As $\arctan(x)$ has slope between $-1$ and $1$, by Eq. 95 we have
@@
Recall $|b|=|f|$ takes values in $[0,1-\eta]$. By the mean value theorem, the Lipschitz constants of the functions $\sqrt{1-x^{2}}$ and $\log(1-x^{2})$ on the interval $[0,1-\eta]$ are at most $\delta^{-1}$ and $2\delta^{-2}$, respectively. Thus we obtain
@@ 5. Complexity analysis of Riemann-Hilbert-Weiss algorithm
Henceforth, we fix $d\in\mathbb{N}$, $0<\eta<\frac{1}{2}$, and $b$ to be a Laurent polynomial of degree $d$ satisfying $b(z)=b(z^{-1})$ and $\left\lVert b\right\rVert_{\infty}\leq 1-\eta$, as in the assumption of the theorem. Also fix $0\leq k\leq d$, $10<\varepsilon<$ 1 and $N$ an even integer as in Theorem 8.
@@ 5.1. Errors analysis for the output $\hat{c}_{j}$ of the Weiss algorithm
is equal to $N$ if $k$ is of the form $j+Nk^{\prime}$ with $k^{\prime}\in\mathbb{Z}$ and equal to 00 otherwise, we obtain
@@
Spliting the sum into positive $k$ and negative $k$ and using geometric decay of the summands, we obtain
@@
where we used that on $\mathbb{T}$ we have that $b$ is bounded above by $1-\eta$ and, using $10<\eta<$ 1,
@@
$\left\lvert b(z)a^{-1}(z)\right\rvert=\left\lvert b(z)\right\rvert\left\lvert a(z)^{*}+b^{*}(z)b(z)a^{-1}(z)\right\rvert\leq r^{d}(1-\eta)(1+r^{d}(1-\eta)\left\lvert b(z)a^{-1}(z)\right\rvert)$
@@
The proposition now follows from Eq. 116 and Eq. 118 and $10<\eta<$ 1. ∎
@@
The following theorem from [28] is our main tool for estimating with the errors coming from the FFT.
Theorem A.
([28]) Let $u$ be an analytic function in the annulus $r^{-1}\leq\left\lvert z\right\rvert\leq r$ for some $r>1$. If $\left\lvert u(z)\right\rvert\leq M$, then for any $K\geq 1$ and any $j$,
(129) $\left\lvert\mathcal{F}(u)(j)\right\rvert\leq Mr^{-\left\lvert j\right\rvert},\quad\left\lvert\mathcal{F}(u)(j)-\mathcal{F}^{K}(u)(j)\right\rvert\leq\frac{M\left(r^{j}+r^{-j}\right)}{r^{K}-1}\,.$
We will apply this theorem to $\frac{b}{a}$ and to $\log(1-|b|^{2})$ according to the two FFTs appearing in Algorithm 1.
@@ 5.2. Error analysis for solving the linear system
$=\left\|\frac{b(z)}{a(z)}\sum_{j=-(d-k)}^{0}x_{j}z^{j}\right\|_{2}\leq\left\lVert\frac{b}{a}\right\rVert_{\infty}\left\lVert\sum_{j=-(d-k)}^{0}x_{j}z^{j}\right\rVert_{2}=\left\lVert\frac{b}{a}\right\rVert_{\infty}\cdot\left\lVert x\right\rVert\,.$
@@
Note that $\hat{r}_{\ell}=\mathcal{F}^{N}\left(\log(1-\left\lvert b(z)\right\rvert^{2})\right)(\ell)$. So $\hat{r}_{\ell}\in\mathbb{R}$ for any $\ell$. Together with that the Fourier coefficients of $b$ are all pure imaginary, we know that the Fourier coefficients of $\hat{u}$ are all pure imaginary. Due to the relation [29, Equations (2.2) and (2.8)], we know that for any $k\leq j\leq d$,
@@ 5.3. Proof of Theorem 8
$|a_{k,0}-\hat{a}_{k,0}|=\left|\begin{pmatrix}\mathbf{e}_{0}^{\top},\mathbf{0}^{\top}\end{pmatrix}\left[(\operatorname{Id}+M_{k})^{-1}-(\operatorname{Id}+\hat{M}_{k})^{-1}\right]\begin{pmatrix}\mathbf{e}_{0}\\ \mathbf{0}\end{pmatrix}\right|$
$=\left|\begin{pmatrix}\mathbf{e}_{0}^{\top},\mathbf{0}^{\top}\end{pmatrix}(\operatorname{Id}+M_{k})^{-1}(\hat{M}_{k}-M_{k})(\operatorname{Id}+\hat{M}_{k})^{-1}\begin{pmatrix}\mathbf{e}_{0}\\ \mathbf{0}\end{pmatrix}\right|$
@@
$\left\|\left(\operatorname{Id}+M_{k}\right)^{-1}\right\|\leq 1\text{ and }\left\lVert\left(\operatorname{Id}+\hat{M}_{k}\right)^{-1}\right\rVert\leq 1\,.$
@@
$b_{k,0}:=\begin{pmatrix}\mathbf{0}^{\top},\mathbf{e}_{0}^{\top}\end{pmatrix}(\operatorname{Id}+M_{k})^{-1}\begin{pmatrix}\mathbf{e}_{0}\\ \mathbf{0}\end{pmatrix},\text{ and }\hat{b}_{k,0}:=\begin{pmatrix}\mathbf{0}^{\top},\mathbf{e}_{0}^{\top}\end{pmatrix}(\operatorname{Id}+\hat{M}_{k})^{-1}\begin{pmatrix}\mathbf{e}_{0}\\ \mathbf{0}\end{pmatrix}\,.$
@@ 5.4. Computational cost
FFT requires $\mathcal{O}(N\log(N))$ operations, which is at most $\mathcal{O}(\frac{d}{\eta}\log(\frac{d}{\eta\varepsilon}))$ operations. The cost for solving a $d\times d$ linear system is $\mathcal{O}(d^{3})$. The overall computational cost of using Algorithm 1 to determine a single phase factor is, then, $\mathcal{O}(d^{3}+\frac{d}{\eta}\log(\frac{d}{\eta\varepsilon}))$. To compute all the phase factors associated to a polynomial $f$ of degree $2d$, the FFT requires still the same number of operations, but we now solve $(d+1)$ many linear systems to compute phase factors, which then takes $\mathcal{O}(d^{4})$ many operations. Thus we are left with a total operational cost of $\mathcal{O}(d^{4}+\frac{d}{\eta}\log(\frac{d}{\eta\varepsilon}))$ to compute all the phase factors of $f$. This completes the analysis of the computational cost as outlined in Theorem 2.
@@ 5.5. Bit requirement
In the Weiss algorithm, we need to apply two FFTs to obtain the Fourier coefficients of $R$ and $b/a$. Due to Parsevel theorem for Discrete Fourier Transform, the 2-norm of the vector of exact FFT coefficients for $R$ is bounded by
@@
B indicates the total absolute error of the Weiss algorithm is
@@
There exists an identity
$(A+E)^{-1}-A^{-1}=(A+E)^{-1}EA^{-1}$
if both $A+E$ and $A$ are nonsingular. We assume that $u$ is sufficiently small and exploit the fact that $\left\lVert\left(1-(\Xi_{k})^{2}\right)^{-1}\right\rVert<1$ for any $k$. Then the forward error of solving this single linear system Eq. 137 is still
@@
where $a_{k,0}$ and $b_{k,0}$ are the first entries of $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$. The previous discussion implies that the error in computing $a_{k,0}$ and $b_{k,0}$ is $\mathcal{O}\left(\frac{d^{2}u}{\eta}\right)$. Mimicking the analysis from Eq. 143, the error of computing $\psi_{k}$ is
@@
$\mathcal{O}\left(\log(\log(N))+\log(d)+\log\left(\frac{1}{\eta}\right)+\log\left(\frac{1}{\varepsilon}\right)\right)\,.$
Plugging in the estimate for $N$ from Theorem 8, we get that the bit required by Algorithm 1 is $\widetilde{\mathcal{O}}(\log(\frac{d}{\varepsilon\eta}))$, which completes the proof of the bit requirement part of Theorem 2.

### v3

@@ 4.2. Riemann-Hilbert factorization: Proof of Theorem 5
where $(a_{-},b_{-})$ is the NLFT of a sequence supported on $(-\infty,-1]$ and $(a_{+},b_{+})$ is the NLFT of a sequence supported on $0,\infty)$, and then it suffices to define
@@
$\lim\limits_{\eta\to 0}\langle A,1\rangle=A(\infty)\,.$
@@ 4.3. Lipschitz estimate
[…]
@@
$\langle\begin{pmatrix}C\\ D\end{pmatrix},\left(\operatorname{Id}+M\right)^{-1}\left(M^{\prime}-M\right)\left(\operatorname{Id}+M^{\prime}\right)^{-1}\begin{pmatrix}1\\ 0\end{pmatrix}\rangle\,$.
@@
Because $\begin{pmatrix}C\\ D\end{pmatrix}\in\mathcal{F}$, then $(\operatorname{Id}-M)^{-1}\begin{pmatrix}C\\ D\end{pmatrix}\in\mathcal{D}\cap\mathcal{D}^{\prime}$ and so by Eq. 60 we may write this last inner product as
@@
$\langle\begin{pmatrix}0&(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})^{*}\\ -(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})&0\end{pmatrix}\left(\operatorname{Id}-M\right)^{-1}\begin{pmatrix}C\\ D\end{pmatrix},\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\rangle\,$,
@@
$\langle\left(\operatorname{Id}-M\right)^{-1}\begin{pmatrix}C\\ D\end{pmatrix},\begin{pmatrix}0&-(\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})^{*}\\ (\frac{b}{a}-\frac{b^{\prime}}{a^{\prime}})&0\end{pmatrix}\begin{pmatrix}A^{\prime}\\ B^{\prime}\end{pmatrix}\rangle\,$,
@@
uniformly in $n$. Without loss of generality, take $n$ to be 0. By [1, (6.13)] and then Eq. 80, we may write
@@
(87) $A(\infty)^{-1}\left\|B-B^{\prime}\right\|_{L^{2}(\mathbb{T})}+A(\infty)^{-1}A^{\prime}(\infty)^{-\frac{1}{2}}\left|\frac{B^{\prime}(0)}{(A^{\prime})(\infty)^{\frac{1}{2}}}\right|\left\|A-A^{\prime}\right\|_{L^{2}(\mathbb{T})}\,.$
@@ 4.4. Plancherel equality and maximal solution
We remark that the particular choice of $a(z)$ given by Theorem 4 actually corresponds to the maximal solution, which is a special class of solutions with symmetric phase-factor proposed in [29] and enjoys many desirable properties. To reveal the connection between the function $a(z)$ constructed by Theorem 4 and the maximal solution, we first review the construction of the $SU(2)$ matrix corresponding to the maximal solution. Without loss of generality, let $f(x)$ be a real even target polynomial of degree $2d$. Then we factorize
@@
is a polynomial and is in $H^{2}(\mathbb{D})$, and it satisfies $\left\lvert\widetilde{a}(z)\right\rvert^{2}=\mathfrak{F}(z)$. Following [29, Theorem 4], we construct
@@
which is the unitary matrix associated with the maximal solution. It is exactly the unitary matrix $U_{d}(x,\Psi)$ associated with $(a(z),b(z))$ by exploiting $a(z^{2})=z^{-4d}\widetilde{a}(z)$ and verifying Eq. 27. Note in particular, that the polynomial $a^{*}$ has no roots in $\overline{\mathbb{D}}$ and therefore must be an outer function. As outer functions are determined by their absolute value on $\mathbb{T}$ by our definition, $a^{*}$ must be the same outer function constructed in the Weiss algorithm. Hence, our algorithm constructs the maximal solution with theoretical guarantees.
@@
If $(a,b)$ is the NLFT of a sequence $F\in\ell^{2}(\mathbb{Z})$, then by the inner-outer factorization theorem for $H^{\infty}(\mathbb{D})$ functions [11, Corollary 5.6, Chapter II], we may write $a^{*}=IO$ where $I$ and $O$ are inner and outer functions on $\mathbb{T}$, respectively. By Eq. 22 we may write
@@
where the last equality follows from Eq. 91. Thus as in [28, (3.1) and the Remark on p.16], we have the nonlinear Plancherel identity
@@ 4.5. Proof of Theorem 1
A careful reading of this section up till now, along with the existence and uniqueness proofs above yields the following result, which is a more comprehensive version of Lemma 6.
@@
As $\arctan(x)$ has slope between 1-1 and $1$, by Eq. 95 we have
@@
Recall $|b|=|f|$ takes values in $0,1-\eta]$. By the mean value theorem, the Lipschitz constants of the functions $\sqrt{1-x^{2}}$ and $\log(1-x^{2})$ on the interval $0,1-\eta]$ are at most $\delta^{-1}$ and $2\delta^{-2}$, respectively. Thus we obtain
@@ 5. Complexity analysis of Riemann-Hilbert-Weiss algorithm
Henceforth, we fix $d\in\mathbb{N}$, $0<\eta<\frac{1}{2}$, and $b$ to be a Laurent polynomial of degree $d$ satisfying $b(z)=b(z^{-1})$ and $\left\lVert b\right\rVert_{\infty}\leq 1-\eta$, as in the assumption of the theorem. Also fix $0\leq k\leq d$, $0<\varepsilon<1$ and $N$ an even integer as in Theorem 8.
@@ 5.1. Errors analysis for the output $\hat{c}_{j}$ of the Weiss algorithm
is equal to $N$ if $k$ is of the form $j+Nk^{\prime}$ with $k^{\prime}\in\mathbb{Z}$ and equal to 0 otherwise, we obtain
@@
Splitting the sum into positive $k$ and negative $k$ and using geometric decay of the summands, we obtain
@@
where we used that on $\mathbb{T}$ we have that $b$ is bounded above by $1-\eta$ and, using $0<\eta<1$,
@@
[…]
@@
The proposition now follows from Eq. 116 and Eq. 118 and $0<\eta<1$. ∎
@@ 5.2. Error analysis for solving the linear system
$\left\|\frac{b(z)}{a(z)}\sum_{j=-(d-k)}^{0}x_{j}z^{j}\right\|_{2}\leq\left\lVert\frac{b}{a}\right\rVert_{\infty}\left\lVert\sum_{j=-(d-k)}^{0}x_{j}z^{j}\right\rVert_{2}=\left\lVert\frac{b}{a}\right\rVert_{\infty}\cdot\left\lVert x\right\rVert\,.$
@@
Note that $\hat{r}_{\ell}=\mathcal{F}^{N}\left(\log(1-\left\lvert b(z)\right\rvert^{2})\right)(\ell)$. So $\hat{r}_{\ell}\in\mathbb{R}$ for any $\ell$. Together with that the Fourier coefficients of $b$ are all pure imaginary, we know that the Fourier coefficients of $\hat{u}$ are all pure imaginary. Due to the relation [27, Equations (2.2) and (2.8)], we know that for any $k\leq j\leq d$,
@@ 5.3. Proof of Theorem 8
[…]
@@
[…]
@@
[…]
@@ 5.4. Computational cost
The FFT for a sequence of length $N$ requires $\mathcal{O}(N\log(N))$ operations. From Theorem 8, we set $N=\mathcal{O}(\frac{d}{\eta}\log(\frac{d}{\eta\varepsilon}))$, and the number of operations of FFT is $\mathcal{O}\left(\frac{d}{\eta}\log(\frac{d}{\eta\varepsilon})\left(\log(\frac{d}{\eta})+\operatorname{loglog}(\frac{d}{\eta\varepsilon})\right)\right)=\mathcal{O}\left(\frac{d}{\eta}\log^{2}(\frac{d}{\eta\varepsilon})\right)$. The cost for solving a $d\times d$ linear system is $\mathcal{O}(d^{3})$. The overall computational cost of using Algorithm 1 to determine a single phase factor is, then $\mathcal{O}(d^{3}+\frac{d}{\eta}\log^{2}(\frac{d}{\eta\varepsilon}))$. To compute all the phase factors associated to a polynomial $f$ of degree $2d$, the FFT requires still the same number of operations, but we now solve $(d+1)$ many linear systems to compute phase factors, which then takes $\mathcal{O}(d^{4})$ many operations. Thus we are left with a total operational cost of $\mathcal{O}(d^{4}+\frac{d}{\eta}\log^{2}(\frac{d}{\eta\varepsilon}))$ to compute all the phase factors of $f$. This completes the analysis of the computational cost as outlined in Theorem 2.
@@ 5.5. Bit requirement
In the Weiss algorithm, we need to apply two FFTs to obtain the Fourier coefficients of $R$ and $b/a$. Due to the Parseval theorem for Discrete Fourier Transform, the 2-norm of the vector of exact FFT coefficients for $R$ is bounded by
@@
A indicates the total absolute error of the Weiss algorithm is
@@
The following identity holds if both $A$ and $A+E$ are nonsingular:
$(A+E)^{-1}-A^{-1}=(A+E)^{-1}EA^{-1}.$
We assume that $u$ is sufficiently small and make use of the bound $\left\lVert(1-(\Xi_{k})^{2})^{-1}\right\rVert<1$ for all $k$. Then the forward error of solving this single linear system Eq. 136 is still
@@
where $a_{k,0}$ and $b_{k,0}$ are the first entries of $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$. The previous discussion implies that the error in computing $a_{k,0}$ and $b_{k,0}$ is $\mathcal{O}\left(\frac{d^{2}u}{\eta}\right)$. Similar to the analysis in Eq. 142, the error of computing $\psi_{k}$ is
@@
$\mathcal{O}\left(\operatorname{loglog}(N)+\log(d)+\log\left(\frac{1}{\eta}\right)+\log\left(\frac{1}{\varepsilon}\right)\right)\,.$
Plugging in the estimate for $N$ from Theorem 8, we find that the bit required by Algorithm 1 is $\mathcal{O}(\log(\frac{d}{\varepsilon\eta}))$, which completes the proof of the bit requirement part of Theorem 2.